OpenStudy (calculusxy):

A line goes through the points (2, 5) and (6, −1). Let P be the point on this line that is closest to the origin. Calculate the coordinates of P.

OpenStudy (calculusxy):

@jhonyy9

OpenStudy (eliesaab):

Find first the equation of the line going thru the 2 points

OpenStudy (eliesaab):

You will find it \[ y=8-\frac{3 x}{2} \]

OpenStudy (eliesaab):

Write it as \[ 3 x + 2y -16=0\]

OpenStudy (mhchen):

Second, find a line that's perpendicular to it, that also goes through the origin.

OpenStudy (eliesaab):

Use now the distance formula from a point\( (m,n)\) to a line a x + by + c=0

OpenStudy (eliesaab):

\[ q=\frac {|a m + b n + c|} {\sqrt{a^2+b^2} } \]

OpenStudy (calculusxy):

Here's what I thought... It seems to be like the x-intercept or the y-intercept might be the closer points to the origin. I calculated the x-int and the y-int and saw that (4.5,0) <- is the closer one

OpenStudy (eliesaab):

replace m=0, n=0, a=3, b=2, c=-16 in the above to get your answer

OpenStudy (calculusxy):

But I am not so sure whether that is exactly correct or not

OpenStudy (eliesaab):

You can also do it by the perpendicular line

OpenStudy (mhchen):

LIke I said ^^ THe shortest distance is basically a radius, and the line is a tangent.

OpenStudy (calculusxy):

I searched through some websites and they said that if I find the perpendicular line that passes through the origin and the point of intersection of the original line and the perpendicular line will give me the point. Is it correct to assume that?

OpenStudy (eliesaab):

\[ \frac {16}{ \sqrt{3^2+2^2}}=\frac {16}{\sqrt{13}} \]

OpenStudy (calculusxy):

I just tried out the thing that I just said recently and my graphing calculator says that the point of intersection is (3.6,1.8)

OpenStudy (mhchen):

Yes, that's what I said in my first post.

OpenStudy (calculusxy):

That's correct right--the coordinates that I just said?

OpenStudy (eliesaab):

The equation of the perpendicular line is y= 2 x/3 in intersects the original line at the point (48/13,32/13)

OpenStudy (eliesaab):

Find the distance from that point to the origin, you find my answer by the firstmethod

OpenStudy (eliesaab):

\[ \sqrt{\left(\frac{32}{13}\right)^2+\left(\frac{48}{13}\right)^2}=\frac{16}{\sqrt{13}} \]

OpenStudy (calculusxy):

The equation of the original line I got is y=-2x+9 The perpendicular line that would cross through the origin would be y = 1/2x

OpenStudy (eliesaab):

No the original equation is y =-3 x/2 +8

OpenStudy (eliesaab):

Recheck your computation

OpenStudy (calculusxy):

got it

OpenStudy (calculusxy):

So now I got approximately (3.69, 2.47)

OpenStudy (calculusxy):

Which I think is about what you got as well

OpenStudy (eliesaab):

Yes

OpenStudy (eliesaab):

Use the distance formula from (0,0) to (48/13,32/13) to get your answer